# Hidden Metallic Iron in Amorphous Silicate Dust? Insights from Condensation Experiments and Mid-Infrared Spectroscopy

### v1

@@ Abstract
Context. Amorphous silicate dust is a major component in the interstellar and circumstellar dust formed in the outflow of asymptotic giant branch (AGB) stars. Although iron depletion is observed in the interstellar medium (ISM), the exact form and fraction of iron in solid remains under debate. In particular, it is unclear whether amorphous silicate dust around AGB stars contains metallic iron.
Aims. We aimed to provide optical constants of amorphous silicate nanoparticles and examine the effects of metallic iron on their spectral features to better constrain the dust properties by producing amorphous silicate nanoparticles with and without metallic cores.
Methods. We performed condensation experiments using an induction thermal plasma system to produce dust analogues of the CI chondritic composition in the Mg–Ca–Na–Al–Si–Fe–Ni–O and Mg–Ca–Na–Al–Si–O systems. We measured the absorbance and reflectance of the samples, observed the structure of the products, and determined the optical constants.
Results. Two types of amorphous silicate nanoparticles () with nearly CI chondritic composition were produced: one contained kamacite (Fe0.9Ni0.1) cores with a diameter ratio ranging 0–0.87 (average 0.50), and the other was iron-free homogeneous amorphous silicate. The amorphous silicates of the CI chondritic composition with various sized metallic cores may be prevalent in circumstellar and interstellar dust.
@@
Infrared spectroscopic observations have revealed that amorphous silicate dust is a dominant refractory dust in circumstellar and interstellar environments. At least 97.8% by mass of silicate dust is amorphous in the diffuse interstellar medium (ISM) (Kemper2004; Kemper2005). Amorphous silicate ubiquitously exists in protoplanetary disks and interplanetary dust particles as well (Henning2010).
@@
Clear evidence of the amorphous silicate formation was observed in outflows of asymptotic giant branch (AGB) stars, evolved stars with low to intermediate mass (). Mid-infrared (MIR) spectra of oxygen-rich AGB stars show amorphous silicate features at around 10 and , attributed to Si–O stretching and Si–O–Si bending vibration, respectively. AGB stars significantly contribute to ISM dust in the Milky Way and their dust injection rate is estimated to be , which is corresponding to nearly 67 % of the total dust injection in the galaxy (Tielens2018). For silicate dust, O-rich AGB stars are estimated to account for 29% of the total silicate dust injection with dust injection rate of (Tielens2005).
@@
Dust grains drive the acceleration of stellar wind by radiation pressure and are ejected into the ISM (Hofner2018). They regulate the thermal structure of circumstellar environments and ISM by absorbing UV radiation and emitting infrared radiation. Dust also significantly affects material evolution by providing surfaces for chemical reactions and acting as catalysts and reactants in ISM and molecular clouds (Cabedo2024; Reboussin2014). Therefore the physical and chemical properties of dust are important to evaluate the functions of dust in the interstellar and circumstellar phenomena.
@@
However, these dust properties around AGB stars are not fully constrained. MIR spectra depend on the optical constants of dust, which change with chemical compositions and structures (e.g., Speck2011; Speck2015), as well as physical parameters such as particle size (Jaeger1994; Dorschner1995), temperature (Zeidler2013), and grain shapes (Fabian2001; Min2005; Min2006). Observations of dust around oxygen-rich AGB stars provide limited diagnostic information because it is challenging to distinguish individual effects of chemical composition, grain size, dust temperature, and shape from broad spectral features of amorphous silicate.
@@
Moreover, it has been a problem where and how iron exists in circumstellar and interstellar environments. Iron is one of the most abundant cationic elements following magnesium and silicon. Thermodynamical equilibrium calculations predict that stable phase of Fe is metallic iron assuming the solar composition at high temperatures, although it can form FeS at lower temperatures or Fe-bearing silicate/silicide at more oxidizing/reductive conditions. Theoretical studies suggest that iron tends to initially condense as a metal rather than form silicates in the circumstellar outflow (Nuth1990; Gail1999; Lodders1999). Even if Fe forms metallic grains, it is crucial to determine whether metallic iron exists inside or separately from the amorphous silicate to understand chemical reactions on the dust surfaces in the ISM. The catalytic function of iron changes depending on its chemical state (whether it is metal or oxides), and on its spatial distribution (whether metal iron exists inside or on the surface of a silicate particle).
@@
Glass with embedded metal and sulfides (GEMS) found in chondritic porous interplanetary dust particles (CP-IDPs) are amorphous silicate grains (Keller2011), including metallic iron nanoparticles inside and sulfide particles on the surface (Matsuno2022). Isotopic analysis of GEMS and experiments forming GEMS analogues suggest part of GEMS may originate from AGB stars (Keller2011; Kim2021; Matsuno2021; Enju2022). Bradley1999 showed that GEMS grains exhibit infrared spectral features similar to those of circumstellar and interstellar amorphous silicates. This spectral match does not necessarily mean that circumstellar dust also has metallic particles. Metallic iron is featureless in the MIR region, and therefore, its effects have been discussed only in terms of near-infrared absorption and opacity enhancement.
@@
Since the 1980s, emission and absorption of circumstellar and interstellar dust have been modeled as hypothetical optical constants referred to as “astronomical silicates” (Draine1984; Ossenkopf1992), which are arbitrarily synthesized by combining the observational data of dust and laboratory measurement. In deriving astronomical silicates, certain optical data of materials containing metallic inclusions are used, such as “dirty silicate” (Jones1976; Rogers1983), in the case of Draine1984, and laboratory silicates [(Mg, Fe)SiO3] with Fe and Fe3O4 inclusions in the case of Ossenkopf1992. However, these optical constants do not represent real solid materials and, therefore, cannot be used to discuss the chemical compositions and structures of dust.
@@
In laboratories, experiments have been carried out to produce amorphous silicate dust analogues with various techniques, such as the sol-gel method (Jaeger2003), the so-called “smoke” method (Nuth2002; Rietmeijer2002), and melt-quenching (Dorschner1995; Mutschke1998). Most of these experiments assume that circumstellar dust is pure amorphous silicate without metallic iron inclusions. On the other hand, Speck2015 proposed the combination of iron-free amorphous silicate with chondritic composition (Na0.11Ca0.12Mg1.86)(Al0.18Si1.85)O6 and metallic iron as an alternative to astronomical silicates, assuming that amorphous silicate grains and metallic iron grains exist separately.
@@
Equilibrium condensation temperature of metallic iron is similar to those of crystalline Mg-silicates (forsterite and enstatite), but kinetically, condensation of metallic iron may delay due to its high surface energy and high efficiency of absorbing stellar light (Kozasa1987; Woitke2006). The abundant presence of amorphous silicates, which are metastable silicate phases, suggests that circumstellar dust formation generally occurs under unequilibrium conditions, and the condensation sequence of amorphous silicates and metallic iron is further unclear. Although Gail1999 estimated that metallic iron first condenses on the surface of silicate, whether the iron exists as separate particles or as inclusions inside the silicate grains is not determined. Kemper2002 argued that non-spherical metallic iron grains explain the observation of AGB star with an optically thick circumstellar dust shell. For the ISM, it is estimated that 70% of the depleted iron in the ISM exists as metallic particles within amorphous silicate by numerical simulations (Zhukovska2018). While Jaeger2003 theoretically showed IR spectral features of dust consisting of amorphous silicate shell and metallic iron core, the effects of metallic Fe–Ni particles on the infrared spectra of amorphous silicate have not been examined experimentally.
@@
Laboratory studies using induction thermal plasma (ITP) systems (Kim2021; Matsuno2021; Enju2022) produced amorphous silicate grains containing metallic iron cores (GEMS-like nanoparticles). The ITP systems can produce nanoparticles for a wider range of chemical compositions than the melt-quenching method by a high temperature () plasma flame to vaporize refractory materials and a rapid cooling condition (). However, the optical constants of amorphous silicate nanoparticles produced with ITP systems, including these GEMS-like nanoparticles, have not been determined. This is because quantitative measurement of reflectance spectra requires polished samples with smooth and flat surfaces, and it is challenging to derive reliable optical constants of nanopowders.
@@
The objective of this study is to present the optical constants of amorphous silicate nanoparticles with the CI chondritic composition, which can be applied to constrain the chemical composition of dust around other AGB stars and to investigate spectral changes with the presence of metallic iron inclusions within amorphous silicate grains. We performed condensation experiments in the Mg–Ca–Na–Al–Si–Fe–Ni–O and Mg–Ca–Na–Al–Si–O systems using the ITP system to produce amorphous silicate nanoparticles with metallic Fe-Ni-containing CI chondritic composition and Fe-Ni-free CI chondritic composition. We determined the optical constants of nanoparticle products by measuring absorption and reflectance spectra and fitting them using the Lorentz oscillator model. With the derived optical constants, we discussed the spectral changes caused by the presence or absence of metallic iron and the possibility that dust around AGB stars has metallic iron particles.
@@ 2.1 Condensation experiments
Experiments were conducted in an induction thermal plasma (ITP) system (JEOL TP-40020NPS, The University of Tokyo; Kim2021) consisting of a plasma torch, chamber, and powder feeder (JEOL TP-99010FDR). High-temperature plasma flame () was generated with the fixed input power of . Ar gas with a flow rate of was used as a main plasma-forming gas. The plasma-forming gas was injected vertically outside the quartz tube and generated a long plasma (radial) flame. The chamber pressure was maintained at during the experiments. The starting material was injected into a plasma flame along with an Ar carrier gas at a flow rate of and vaporized with the plasma flame. The gas was quenched by water running outside the chamber () to condense nanoparticles. The feeding rates of the starting materials were calculated from the total mass of the products and the experimental duration of 15 minutes. The experimental conditions are shown in Table 1.
@@
The starting material was prepared by grinding oxides and metallic power reagents: MgO (¿99.99% purity, ; Kojundo Chemical Lab. Co., Ltd.), CaCO3 (¿99% purity, ; Kojundo Chemical Lab. Co., Ltd.), Na2SiO3 (¿99% purity, prepared by grounding coarse grains; Kojundo Chemical Lab. Co., Ltd.), Al (¿99.9% purity, ; Kojundo Chemical Lab. Co., Ltd.), Fe (¿99.9% purity, ; Kojundo Chemical Lab. Co., Ltd.), Ni (¿99% purity, ; Kojundo Chemical Lab. Co., Ltd.), SiO2 ( purity, ; Kojundo Chemical Lab. Co., Ltd.), Si (¿% purity, ; Kojundo Chemical Lab. Co., Ltd.). These reagents were mixed at the CI chondritic composition (Lodders2009) in the Mg–Ca–Na–Al–Si–Fe–Ni–O and Mg–Ca–Na–Al–Si–O systems (Table 2).
@@
The redox condition in the ITP system was estimated by the total oxygen abundance in the system relative to the oxygen abundance to convert cations other than Fe and Ni into oxides, (Enju2022):
Where is the feeding rate of the starting materials, is the oxygen inflow rate into the ITP system, which is estimated to be by Enju2022, is the oxygen abundance of the starting material and Ao is the oxygen abundance of the starting material, assuming that all Si and Al are fully oxidized. Fe condenses as a metal under reducing conditions at , and is incorporated into silicates as a cation under oxidizing conditions at . In the Fe-Ni-containing system, we aimed to obtain amorphous silicates with Fe-Ni metal inclusions by controlling with the range of 0.93-0.98 with setting the SiO2 / (Si + SiO2) ratio to 0.95 in the starting materials. In the Fe-Ni-free system, was set to be 9.35-9.40. All Si was introduced as Si powders instead of SiO2, and O2 plasma-forming gas was added at a flow rate of to prevent the formation of any metallic particles and silicides.
The condensates were collected from the upper and lower walls of the chamber separately. With a radial flame, the size distribution of the condensates does not differ with the distance from the plasma flame in a chamber. The bulk chemical composition and peak position in absorption spectra were almost identical between the products from the upper and lower walls. As more condensates were collected from the lower wall, we used the lower-wall product for analyses.
@@ 2.2 Analytical methods
The bulk chemical composition of the products was measured using an electron probe microanalyzer (EPMA; JEOL JXA-8530F, the University of Tokyo). The condensates were pressed onto Cu plates at for 1 minute using an oil hydraulic press and subsequently coated with carbon. The pressed samples were analyzed with a defocused electron beam of in diameter at an accelerating voltage of and a beam current of . Approximately 20 points were selected from the homogeneous sample surface regions without large particles ( in diameter), which were supposed to be residues of the starting materials. Albite (NaAlSi3O8), periclase (MgO), Al2O3, wollastonite (CaSiO3), Fe2O3, and NiO were used as standard materials. The measured data were corrected using the ZAF correction.
The structures of the condensates were observed, and the chemical composition of the nanoparticles was measured using transmission electron microscopy (TEM; JEOL JEM-2800, the University of Tokyo) equipped with energy dispersive X-ray spectroscopy (EDS). The condensed particles were mounted onto Cu TEM grids coated with carbon without any dispersant. Dark field scanning transmission electron microscopy (DF-STEM) images and STEM-EDS maps with a resolution of pixels were obtained. To minimize beam damage to the sample, a beam with an accelerating voltage of and a spot size of was used for a one-hour measurement.
The products were analyzed with Fourier transform infrared (FT-IR) spectroscopy to obtain transmittance and reflectance spectra. For measuring transmittance spectra, the condensates were dispersed in KBr powder at a mass ratio of 1:500, evacuated for 10 minutes, and pressed at for 10 minutes into diameter disks using an oil hydraulic press under vacuum. The transmittance spectra of the sample pellets were obtained over the range of with resolution of (JASCO FT/IR-4200, The University of Tokyo).
For reflectance measurements, only the condensed nanopowders were pressed into pellets. In addition to a gold mirror, which is commonly used as a reference for reflectance measurements, we prepared four reference materials to evaluate the scattering effects caused by surface roughness of the pellets made of nanoparticles. The pellets of CI-1 and CI-2 were coated with gold of approximately and thickness using a sputter coater (HITACHI sputter coater E-1030, The University of Tokyo). Although CI-1 and CI-2 pellets were coated with gold simultaneously, a slight difference in their distances from the evaporation source resulted in a thinner Au coating on CI-1 than on CI-2. Reflectance spectra were measured at an incident light angle of over the mid-infrared (MIR) range of with a resolution of (Thermo NICOLET6700, Osaka Sangyo University).

### v2

@@ Abstract
Context. Amorphous silicate dust is a major component in the interstellar and circumstellar dust formed in the outflow of asymptotic giant branch (AGB) stars. Although iron depletion is observed in the interstellar medium, the exact form and fraction of iron in solids remains a matter of debate. In particular, it is unclear whether the amorphous silicate dust around AGB stars contains metallic iron.
Aims. We aimed to provide optical constants of amorphous silicate nanoparticles and examine the effects of metallic iron on their spectral features to better constrain the dust properties. We did this by producing amorphous silicate nanoparticles with and without metallic cores.
Methods. We performed condensation experiments using an induction thermal plasma system to produce dust analogs of the CI chondritic composition in Mg–Ca–Na–Al–Si–Fe–Ni–O and Mg–Ca–Na–Al–Si–O systems. We measured the absorbance and reflectance of the samples, observed the structure of the products, and determined the optical constants.
Results. Two types of amorphous silicate nanoparticles () with nearly CI chondritic compositions were produced: one contained kamacite (Fe0.9Ni0.1) cores with a diameter ratio ranging from 0–0.87 (average 0.50), and the other was iron-free homogeneous amorphous silicate. The amorphous silicates of the CI chondritic composition with various-sized metallic cores may be prevalent in circumstellar and interstellar dust.
@@
Infrared (IR) spectroscopic observations have revealed that amorphous silicate dust is a dominant refractory dust in circumstellar and interstellar environments. At least 97.8% of silicate dust (by mass) is amorphous in the diffuse interstellar medium (ISM; Kemper2004; Kemper2005). Amorphous silicate ubiquitously exists in protoplanetary disks and interplanetary dust particles as well (Henning2010).
@@
Clear evidence of amorphous silicate formation was observed in outflows of asymptotic giant branch (AGB) stars, which are evolved stars with low to intermediate masses (). Mid-infrared (MIR) spectra of oxygen-rich AGB stars show amorphous silicate features at around 10 and , which is attributed to Si–O stretching and Si–O–Si bending vibration, respectively. AGB stars contribute significantly to the ISM dust in the Milky Way; their dust injection rate is estimated to be , which corresponds to nearly 67% of the total dust injection in the Galaxy (Tielens2018). For silicate dust, O-rich AGB stars are estimated to account for 29% of the total silicate dust injection, with a dust injection rate of (Tielens2005).
@@
Dust grains drive the acceleration of stellar wind via radiation pressure and are ejected into the ISM (Hofner2018). They regulate the thermal structure of circumstellar environments and the ISM by absorbing UV radiation and emitting IR radiation. Dust also significantly affects material evolution by providing surfaces for chemical reactions and acting as catalysts and reactants in the ISM and molecular clouds (Cabedo2024; Reboussin2014). Therefore, understanding the physical and chemical properties of dust is important when evaluating the functions of dust in interstellar and circumstellar phenomena.
@@
However, the properties of dust around AGB stars are not fully constrained. The shape of MIR spectra depends on the optical constants of dust, which change with chemical compositions and structures (e.g., Speck2011; Speck2015), as well as physical parameters such as particle size (Jaeger1994; Dorschner1995), temperature (Zeidler2013), and grain shapes (Fabian2001; Min2005; Min2006). Observations of dust around oxygen-rich AGB stars provide limited diagnostic information because it is challenging to distinguish individual effects of chemical composition, grain size, dust temperature, and shape from broad spectral features of amorphous silicate.
@@
Moreover, determining where and how iron exists in circumstellar and interstellar environments has been challenging. Iron is one of the most abundant cationic elements, following magnesium and silicon. Thermodynamic equilibrium calculations predict that the stable phase of Fe is metallic iron, assuming the solar composition at high temperatures, although it can form FeS at lower temperatures or Fe-bearing silicate under more oxidizing conditions, and Fe-bearing silicide under more reductive conditions. Theoretical studies suggest that iron tends to initially condense as a metal rather than form silicates in the circumstellar outflow (Nuth1990; Gail1999; Lodders1999). Even if Fe forms metallic grains, it is crucial to determine whether metallic iron exists inside or separately from amorphous silicate to understand chemical reactions on the dust surfaces in the ISM. The catalytic function of iron changes depending on its chemical state (whether it is metal or an oxide), and on its spatial distribution (whether metal iron exists inside or on the surface of a silicate particle).
@@
Glass with embedded metal and sulfides (GEMS) found in chondritic porous interplanetary dust particles are amorphous silicate grains (Keller2011) that contain metallic iron nanoparticles inside and sulfide particles on the surface (Matsuno2022). Isotopic analysis of GEMS and experiments forming GEMS analogs suggest that some GEMS originates from AGB stars (Keller2011; Kim2021; Matsuno2021; Enju2022). Bradley1999 showed that GEMS grains exhibit IR spectral features similar to those of circumstellar and interstellar amorphous silicates. This spectral match does not necessarily mean that circumstellar dust also has metallic particles. Metallic iron is featureless in the MIR region, and therefore, its effects have been discussed only in terms of near-infrared (NIR) absorption and opacity enhancement.
@@
Since the 1980s, emission and absorption of circumstellar and interstellar dust have been modeled as hypothetical optical constants referred to as “astronomical silicates” (Draine1984; Ossenkopf1992), which are arbitrarily synthesized by combining the observational data of dust and laboratory measurements. In deriving astronomical silicates, certain optical data of materials containing metallic inclusions are used, such as “dirty silicate” (Jones1976; Rogers1983), in the case of Draine1984, and laboratory silicates [(Mg, Fe)SiO3] with Fe and Fe3O4 inclusions in the case of Ossenkopf1992. However, these optical constants do not represent real solid materials and, therefore, cannot be used to discuss the chemical compositions and structures of dust.
@@
In laboratories, experiments have been carried out to produce amorphous silicate dust analogs with various techniques, such as the sol-gel method (Jaeger2003), the so-called “smoke” method (Nuth2002; Rietmeijer2002), and melt-quenching (Dorschner1995; Mutschke1998). Most of these experiments assume that circumstellar dust is pure amorphous silicate without metallic iron inclusions. On the other hand, Speck2015 proposed the combination of iron-free amorphous silicate with a chondritic composition, (Na0.11Ca0.12Mg1.86)(Al0.18Si1.85)O6, and metallic iron as an alternative to astronomical silicates, assuming that amorphous silicate grains and metallic iron grains exist separately.
@@
The equilibrium condensation temperature of metallic iron is similar to those of crystalline Mg-silicates (forsterite and enstatite), but kinetically, the condensation of metallic iron can be delayed due to its high surface energy and high efficiency of absorbing stellar light (Kozasa1987; Woitke2006). The abundant presence of amorphous silicates, which are metastable silicate phases, suggests that circumstellar dust formation generally occurs under nonequilibrium conditions, and the condensation sequence of amorphous silicates and metallic iron is further unclear. Although Gail1999 estimated that metallic iron first condenses on the surface of silicate, whether the iron exists as separate particles or as inclusions inside the silicate grains has not been determined. Kemper2002 argued that nonspherical metallic iron grains explain the observation of an AGB star with an optically thick circumstellar dust shell. In the ISM, an estimated 70% of its depleted iron exists as metallic particles within amorphous silicate, according to numerical simulations (Zhukovska2018). While Jaeger2003 theoretically showed IR spectral features of dust consisting of an amorphous silicate shell and a metallic iron core, the effects of metallic Fe–Ni particles on the IR spectra of amorphous silicate have not been examined experimentally.
@@
Laboratory studies using induction thermal plasma (ITP) systems (Kim2021; Matsuno2021; Enju2022) produced amorphous silicate grains containing metallic iron cores (GEMS-like nanoparticles). The ITP systems can produce nanoparticles for a wider range of chemical compositions than the melt-quenching method by using a high-temperature plasma flame () to vaporize refractory materials and a rapid cooling condition (). However, the optical constants of amorphous silicate nanoparticles produced with ITP systems, including these GEMS-like nanoparticles, have not been determined. This is because the quantitative measurement of reflectance spectra requires polished samples with smooth and flat surfaces, and it is challenging to derive reliable optical constants of nanopowders.
@@
The objective of this study was to determine the optical constants of amorphous silicate nanoparticles with a CI chondritic composition, which can be applied to constrain the chemical composition of dust around other AGB stars and to investigate spectral changes with the presence of metallic iron inclusions within amorphous silicate grains. We performed condensation experiments in Mg–Ca–Na–Al–Si–Fe–Ni–O and Mg–Ca–Na–Al–Si–O systems using the ITP system to produce amorphous silicate nanoparticles with a metallic Fe-Ni-containing CI chondritic composition and a Fe-Ni-free CI chondritic composition. We determined the optical constants of nanoparticle products by measuring absorption and reflectance spectra and fitting them using the Lorentz oscillator model. With the derived optical constants, we discuss the spectral changes caused by the presence or absence of metallic iron and the possibility that dust around AGB stars has metallic iron particles.
@@ 2.1 Condensation experiments
Experiments were conducted in an ITP system (JEOL TP-40020NPS, The University of Tokyo; Kim2021) consisting of a plasma torch, chamber, and powder feeder (JEOL TP-99010FDR). High-temperature plasma flame () was generated with a fixed input power of . Ar gas with a flow rate of was used as a main plasma-forming gas. The plasma-forming gas was injected vertically outside the quartz tube and generated a long plasma (radial) flame. The chamber pressure was maintained at during the experiments. The starting material was injected into a plasma flame along with an Ar carrier gas at a flow rate of and vaporized with the plasma flame. The gas was quenched by water running outside the chamber () to condense nanoparticles. The feeding rates of the starting materials were calculated from the total mass of the products and the experimental duration of 15 minutes. The experimental conditions are shown in Table 1.
@@
The starting material was prepared by grinding oxides and metallic power reagents: MgO (¿99.99% purity, ; Kojundo Chemical Lab. Co., Ltd.), CaCO3 (¿99% purity, ; Kojundo Chemical Lab. Co., Ltd.), Na2SiO3 (¿99% purity, prepared by grounding coarse grains; Kojundo Chemical Lab. Co., Ltd.), Al (¿99.9% purity, ; Kojundo Chemical Lab. Co., Ltd.), Fe (¿99.9% purity, ; Kojundo Chemical Lab. Co., Ltd.), Ni (¿99% purity, ; Kojundo Chemical Lab. Co., Ltd.), SiO2 ( purity, ; Kojundo Chemical Lab. Co., Ltd.), and Si (¿% purity, ; Kojundo Chemical Lab. Co., Ltd.). These reagents were mixed at the CI chondritic composition (Lodders2009) in the Mg–Ca–Na–Al–Si–Fe–Ni–O and Mg–Ca–Na–Al–Si–O systems (Table 2).
@@
The redox condition in the ITP system was estimated by using the total oxygen abundance in the system relative to the oxygen abundance to convert cations other than Fe and Ni into oxides, (Enju2022):
Here is the feeding rate of the starting materials, is the oxygen inflow rate into the ITP system, which is estimated to be by Enju2022, is the oxygen abundance of the starting material, and Ao is the oxygen abundance of the starting material, assuming that all Si and Al are fully oxidized. Fe condenses as a metal under reducing conditions at , and is incorporated into silicates as a cation under oxidizing conditions at . In the Fe-Ni-containing system, we aimed to obtain amorphous silicates with Fe-Ni metal inclusions by controlling with the range of 0.93-0.98, by setting the SiO2 / (Si + SiO2) ratio to 0.95 in the starting materials. In the Fe-Ni-free system, was set to be 9.35-9.40. All Si was introduced as Si powders instead of SiO2, and O2 plasma-forming gas was added at a flow rate of to prevent the formation of any metallic particles and silicides.
The condensates were collected from the upper and lower walls of the chamber separately. With a radial flame, the size distribution of the condensates does not differ with the distance from the plasma flame in a chamber. The bulk chemical composition and peak position in absorption spectra of the products from the upper and lower walls were almost identical. As more condensates were collected from the lower wall, we used the lower-wall product for analyses.
@@ 2.2 Analytical methods
The bulk chemical composition of the products was measured using an electron probe microanalyzer (EPMA; JEOL JXA-8530F, the University of Tokyo). The condensates were pressed onto Cu plates at for 1 minute using an oil hydraulic press and subsequently coated with carbon. The pressed samples were analyzed with a defocused electron beam of in diameter at an accelerating voltage of and a beam current of . Approximately 20 points were selected from the homogeneous sample surface regions without large particles ( in diameter), which were supposed to be residues of the starting materials. Albite (NaAlSi3O8), periclase (MgO), Al2O3, wollastonite (CaSiO3), Fe2O3, and NiO were used as standard materials. The measured data were corrected using the ZAF (atomic number, absorption, and fluorescence) correction.
The structures of the condensates were observed and the chemical composition of the nanoparticles was measured using transmission electron microscopy (TEM; JEOL JEM-2800, the University of Tokyo) equipped with energy dispersive X-ray spectroscopy (EDS). The condensed particles were mounted onto Cu TEM grids coated with carbon without any dispersant. Dark field scanning transmission electron microscopy (STEM) images and STEM-EDS maps with a resolution of pixels were obtained. To minimize beam damage to the sample, a beam with an accelerating voltage of and a spot size of was used for a one-hour measurement.
The products were analyzed with Fourier transform infrared spectroscopy (FT-IR) to obtain transmittance and reflectance spectra. For measuring transmittance spectra, the condensates were dispersed in KBr powder at a mass ratio of 1:500, evacuated for 10 minutes, and pressed at for 10 minutes into diameter disks using an oil hydraulic press under vacuum. The transmittance spectra of the sample pellets were obtained over the range of with a resolution of (JASCO FT/IR-4200, The University of Tokyo).
For reflectance measurements, only the condensed nanopowders were pressed into pellets. In addition to a gold mirror, which is commonly used as a reference for reflectance measurements, we prepared four reference materials to evaluate the scattering effects caused by the surface roughness of the nanoparticle-made pellets. The CI-1 and CI-2 pellets were coated with gold of approximately and thickness using a sputter coater (HITACHI sputter coater E-1030, The University of Tokyo). Although the CI-1 and CI-2 pellets were coated with gold simultaneously, a slight difference in their distances from the evaporation source resulted in a thinner Au coating on CI-1 than on CI-2. Reflectance spectra were measured at an incident light angle of over the MIR range of with a resolution of (Thermo NICOLET6700, Osaka Sangyo University).